# Counterintuitive Potential-Barrier Affinity Effect

### v2

@@ Abstract
Electron accumulation in interatomic regions is a fundamental quantum phenomenon dictating chemical bonding and material properties, yet its origin remains elusive across disciplines[1, 2, 3, 4]. Here, we report a counterintuitive quantum effect—potential-barrier affinity (PBA)—revealed by solving the Schrödinger equation for a crystalline potential. PBA effect drives significant interatomic electron accumulation when electron energy exceeds the barrier maximum. This effect essentially determines interatomic electron density patterns, governing microstructures and properties of condensed matters. Our theory overturns the traditional wisdom that the interstitial electron localization in electride requires potential-well constraints[2, 3, 4] or hybrid orbitals[5], and it serves the fundamental mechanism underlying the formation of conventional chemical bonds. This work delivers a paradigm shift in understanding electron distribution and establishes a theoretical foundation for the microscopic design of material properties.
@@ Results
We subsequently calculate the effective local potential by the Kohn-Sham density functional theory[20] in Figure 2a, along with the valence band maximum (VBM) and conduction band minimum (CBM). Contrary to the conventional expectation of potential wells, we found that electrons localize at the maximum point of the potential ($V_{M}$) within the interstitial region. Furthermore, we observed that the Fermi energy level ($E_{F}=16.0$ eV) is much higher than the maximum point of the potential ($V_{M}=4.4$ eV), indicating that near-free electrons (NFEs) occupy the single-particle states with energy levels ranging from $V_{M}$ to $E_{F}$. The band structure shown in Figure 2b (left) displays two valence electron bands (4 electrons per cell) located within the NFE energy range. This suggests that all valence electrons behave as unbound-state NFEs rather than being confined as interstitial bound electrons. Additionally, the density of states (DOS) in Figure 2b (right) demonstrates that the s, p, and d components are negligible compared to the total DOS within the valence-electron energy range, which is also a distinct feature commonly associated with unbound-state NFEs[21], rather than the multicentered bonding of hybrid orbitals. To further clearly demonstrate the presence of the electron localization within the interstitial region, we plot the Kohn-Sham effective local potential ($V_{eff}^{KS}$), total ($\rho$), bound-state ($\rho_{B}$), and near-free ($\rho_{F}$) electron densities along the $\bar{1}1$] direction of Na-hP4 in Figure 2c. We observed that the NFE bands contribute almost all the interstitial electron density near the maximum potential point, while the electron density tails from the bound-state bands can be ignored. This result directly confirms that interstitial electrons originate from near-free-valence states. We also present the effective ionic potential of Na+-hP4 in Figure 5 of the Method Section by removing the valence electrons. The analysis further confirms that IAEs behave as NFEs with high density accumulation around the potential maximum. Notably, we further found evidence for the existence of NFEs ($E_{F}>V_{M}$) in representative electrides[22, 23, 8, 24, 25, 26, 27], as illustrated in Figure 2d, suggesting that the existence of near-free valence electrons is a common phenomenon among electrides. These findings compel a reevaluation of the current understanding of IAEs in electrides.
@@ Discussion
In summary, we theoretically reveal a universal yet counterintuitive PBA effect by solving the Schrödinger eqution of one-dimensional crystal potential. The PBA effect is characterized by significant electron accumulation within the interatomic region, arising from the smooth connection of wavefunctions between short-wavelength components in atomic region and long-wavelength components across the intetaromic region, amplifying the latter and producing electron accumulation there. We demonstrate that PBA effect becomes most pronounced as the electron energy approaches the barrier maximum and fundamentally governs the resulting electron density distribution in condensed matter, thereby dictating their microstructures and properties. This theory overturns the traditional belief that the formation of IAEs strictly requires potential well constraints or multicentered bonds, and it serves as the fundamental mechanism underlying the formation of conventional chemical bonds, as well as the excited-state electron density distribution. This unexpected discovery represents a fundamental paradigm shift in understanding electron behavior in materials, providing a universal theoretical foundation for the microscopic design of specific material properties.
@@ Methods
We emphasize that in region A, the wavefunction must decay exponentially, as indicated by the Schrödinger equation for a classically forbidden region. This behavior is fundamentally different from the planewave-like character of the wavefunctions in regions B, C, and D. With the eight numerical conditions (six continuity conditions at the points of $x=\{x_{1},x_{2},x_{3}\}$; two periodic boundary conditions at $x=0$ and $x=x_{4}$), the eight coefficients in Eq. (11) can be strictly solved. The potential-barrier affinity effect can occur in region C with planewave-like wavefunctions, as demonstrated in the maintext of our manuscript. We note that systematically adjusting the relative height and spatial separation of the potential barriers in this generalized model may yield deeper insights into the understanding of the formation mechanisms of different bonding types. For the AE KSDFT calculations, we performed the calculations using the full-potential linearised augmented-planewave method[37] by the ELK-v10.6.2 code[38]. $18\times 18\times 18$ $k$-grid meshes are used for the ground-state calculations of the Diamond crystal. The potential and electron density are represented in the dense real-space grids of $120\times 120\times 120$.

### v3

# Potential-Barrier Affinity Effect in Solid Systems

@@ Abstract
Electron accumulation in interatomic regions is a fundamental quantum phenomenon dictating chemical bonding and material properties, yet its origin remains elusive across disciplines[1, 2, 3, 4]. Here, we report a quantum accumulation effect—potential-barrier affinity (PBA)—revealed by solving the Schrödinger equation for a crystalline potential. PBA effect drives significant interatomic electron accumulation when electron energy exceeds the barrier maximum. This effect essentially enhances interatomic electron density, governing microstructures and properties of condensed matter. Our theory overturns the traditional wisdom that the interstitial electron localization in electride requires potential-well constraints[2, 3, 4] or hybrid orbitals[5], and it serves as the fundamental mechanism underlying the formation of conventional solid bonding. This work delivers a paradigm shift in understanding electron distribution and establishes a theoretical foundation for the microscopic design of material properties.
@@ Results
We subsequently calculate the effective local potential by the Kohn-Sham density functional theory[20] in Figure 2a, along with the valence band maximum (VBM) and conduction band minimum (CBM). Contrary to the conventional expectation of potential wells, we found that electrons localize at the maximum point of the potential ($V_{M}$) within the interstitial region. Furthermore, we observed that the Fermi energy level ($E_{F}=16.0$ eV) is much higher than the maximum point of the potential ($V_{M}=4.4$ eV), indicating that near-free electrons (NFEs) occupy the single-particle states with energy levels ranging from $V_{M}$ to $E_{F}$. The band structure shown in Figure 2b (left) displays two valence electron bands (4 electrons per cell) located within the NFE energy range. This suggests that all valence electrons behave as unbound-state NFEs rather than being confined as interstitial bound electrons. While early models interpreted the interstitial-potential well as arising from Pauli exclusion by core electrons[21, 2], a quantitatively accurate description of the Pauli potential remains absent within the fermion quasiparticle framework. We address this point further in the Method section.
Additionally, the density of states (DOS) in Figure 2b (right) demonstrates that the s, p, and d components are negligible compared to the total DOS within the valence-electron energy range, which is also a distinct feature commonly associated with unbound-state NFEs[22], rather than the multicentered bonding of hybrid orbitals. In addition, the concept of hybrid orbital derives from a linear combination of localized atomic orbitals centered on a single atom site[23]. As such, it should be applied with caution when describing unbound states. To further clearly demonstrate the presence of the electron localization within the interstitial region, we plot the Kohn-Sham effective local potential ($V_{eff}^{KS}$), total ($\rho$), bound-state ($\rho_{B}$), and near-free ($\rho_{F}$) electron densities along the $\bar{1}1$] direction of Na-hP4 in Figure 2c. We observed that the NFE bands contribute almost all the interstitial electron density near the maximum potential point, while the electron density tails from the bound-state bands can be ignored. This result directly confirms that interstitial electrons originate from near-free-valence states. We also present the effective ionic potential of Na+-hP4 in Figure 5 of the Method Section by removing the valence electrons. The analysis further confirms that IAEs behave as NFEs with high density accumulation around the potential maximum. Notably, we further found evidence for the existence of NFEs ($E_{F}>V_{M}$) in representative electrides[24, 25, 8, 26, 27, 28, 29], as illustrated in Figure 2d, suggesting that the existence of near-free valence electrons is a common phenomenon among electrides. More evidence of the NFE-density distribution in the interstitial region can be found in Figure 6 and Figure 7 of the Method section. These findings compel a reevaluation of the current understanding of IAEs in electrides.
@@ Discussion
In summary, we theoretically reveal a universal PBA effect by solving the Schrödinger equation of a one-dimensional crystal potential. The PBA effect is characterized by significant electron accumulation within the interatomic region, arising from the smooth connection of wavefunctions between short-wavelength components in the atomic region and long-wavelength components across the interatomic region, amplifying the latter and producing electron accumulation there. We demonstrate that the PBA effect becomes most pronounced as the electron energy approaches the barrier maximum and fundamentally governs the resulting electron density distribution in condensed matter, thereby dictating their microstructures and properties. From the PBA effect, we derive the necessary conditions for the existence of electrides: (i) the Fermi level lies above the maximum of the effective potential barrier; (ii) the barrier maximum is located in a spatially open region far from nuclei; and (iii) an appreciable density of occupied states exists between the Fermi level and the barrier maximum. Such conditions provide a physically transparent guideline for the discovery of an electride. Furthermore, this theory overturns the traditional belief that the formation of IAEs strictly requires potential well constraints or multicentered bonding, and it serves as the fundamental mechanism underlying the formation of conventional solid bonding, defect states, and excited-state electron density distribution. This unexpected discovery represents a fundamental paradigm shift in understanding electron behavior in materials, providing a universal theoretical foundation for the microscopic design of specific material properties.
@@ Methods
Near-free electrons in electrides
Figure 6: The electron density distribution functions for near-free electrons ($\rho_{F}$) and bound-state electrons ($\rho_{B}$), where $\rho_{F}$ and $\rho_{B}$ are calculated according to the unbound-state bands ($V_{M}<\epsilon_{nk}<E_{F}$) and bound-state bands ($\epsilon_{nk}<V_{M}$). The isosurfaces of 0.008, 0.005, 0.008, 0.01, 0.04, 0.1, and 0.05 Å-3 are used for Ca2N, Ba2N, Sr2N, Li5Si, Li6Al, Li8Au, and Li4Rh, respectively.
Figure 7: (a) The electron density distribution functions for near-free electrons ($\rho_{F}$), and (b) the effective potential and electron density calculated along the dashed line of the Ca-Ca pair for Ca24Al28O64, where the isosurface is 0.001 Å-3.
To further establish that IAEs arise from near-free-electron (NFE) behavior, we calculated the partial electron density from band decomposition for the electrides in Fig. 2(d) of the maintext, as well as the representative mayenite [Ca24Al28O64][11]. he results, presented in the Figure 6 and Figure 7, reveal that the NFE density accumulates predominantly in the interstitial region, in agreement with the findings for Na-hP4.
@@
We emphasize that in region A, the wavefunction must decay exponentially, as indicated by the Schrödinger equation for a classically forbidden region. This behavior is fundamentally different from the planewave-like character of the wavefunctions in regions B, C, and D. With the eight numerical conditions (six continuity conditions at the points of $x=\{x_{1},x_{2},x_{3}\}$; two periodic boundary conditions at $x=0$ and $x=x_{4}$), the eight coefficients in Eq. (11) can be strictly solved. The potential-barrier affinity effect can occur in region C with planewave-like wavefunctions, as demonstrated in the maintext of our manuscript. We note that systematically adjusting the relative height and spatial separation of the potential barriers in this generalized model may yield deeper insights into the understanding of the formation mechanisms of different bonding types.
Pauli potential
Figure 9: (a) The interacting fermion systems and reference noninteracting systems. The all-electron KSDFT calculations of effective potentials and densities for (b) Na-hP4 along $1\bar{1}$1] direction, (c) FCC-Al along [111] direction, and (d) Diamond along [111] direction.
Although early theoretical models attributed the interstitial-potential well to Pauli exclusion from core electrons[21, 2], no accurate or quantitative theory exists within the fermionic quasiparticle framework to describe the Pauli potential. Here, to comprehensively analyze the impact of Pauli exclusion in the PBA effect, we mapped the noninteracting fermions (Kohn-Sham system) to the reference noninteracting bosons, where the bosons have the same charge, mass, point-like shape, chemical potential, and density as the electrons[39, 40], as shown in Fig. 9(a). In KSDFT, the electronic total energy of the interacting systems is defined as:
$E[\rho]=T_{s}[\rho]+E_{H}[\rho]+E_{XC}[\rho]+E_{NE}[\rho],$ (12)
where $E_{H}$, $E_{XC}$, and $E_{NE}$ are the Hartree energy, exchange-correlation energy, and nuclei-electron interacting energy, respectively. With the reference system of noninteracting bosons, the noninteracting kinetic energies ($T_{s}$) can be rewritten as[40]:
$T_{s}[\rho]=T_{vW}[\rho]+T_{P}[\rho],$ (13)
where $T_{vW}[\rho]=-\frac{1}{2}\int{\sqrt{\rho({\mathbf{r}})}\nabla^{2}\sqrt{\rho({\mathbf{r}})}}d^{3}{\mathbf{r}}$ is the von Weizsäcker (vW) kinetic energy[41] for the noninteracting bosons occupying the single collective orbital $\phi_{B}({\mathbf{r}})=\sqrt{\rho({\mathbf{r}})}$ (or Madelung wavefunction[42]). The Pauli kinetic energy ($T_{P}$) is defined as the difference between the vW kinetic energy and the exact one due to the Pauli exclusion. Then, we can derive the effective potential by performing the variation on the total energy functional to be stationary with the conservation of electron number:
$\mu=\frac{\delta E[\rho]}{\delta\rho({\mathbf{r}})}=$ $\frac{\delta T_{v}W[\rho]}{\delta\rho({\mathbf{r}})}+\frac{\delta T_{P}[\rho]}{\delta\rho({\mathbf{r}})}+\frac{\delta(E_{H}[\rho]+E_{XC}[\rho]+E_{NE}[\rho])}{\delta\rho({\mathbf{r}})}$
$=$ $V_{vW}[\rho]({\mathbf{r}})+V_{P}[\rho]({\mathbf{r}})+V_{eff}^{KS}[\rho]({\mathbf{r}}),$ (14)
where $V_{vW}[\rho]({\mathbf{r}})=-\frac{1}{2}{\nabla^{2}\sqrt{\rho({\mathbf{r}})}}/{\sqrt{\rho({\mathbf{r}})}}$ and $V_{eff}^{KS}[\rho]({\mathbf{r}})$ are the vW kinetic potential and Kohn-Sham effective potential. Although the exact and analytical Pauli potential ($V_{P}$) is still unknown, we can numerically calculate the effective potential for the reference bosons using the ground-state electron density and chemical potential calculated from KSDFT:
$V_{eff}^{B}[\rho]({\mathbf{r}})=V_{P}[\rho]({\mathbf{r}})+V_{eff}^{KS}[\rho]({\mathbf{r}})=\mu-V_{vW}[\rho]({\mathbf{r}}),$ (15)
where the chemical potential is set as the Fermi energy level of KSDFT in this work ($\mu=E_{F}$). To include the interaction between nuclear and valence electrons, we performed the all-electron (AE) KSDFT calculations using the full-potential linearised augmented-planewave method[43] by the ELK-v10.6.2 code[44]. The AE effective potentials and electron densities are presented in Figure 9(c-d) for Na-hP4 $\bar{1}1$], Al-FCC [111], and Diamond [111] directions. In this framework, the collective quantum state exhibits bound-state behavior under the all-electron effective potential containing Pauli interaction. The results completely overwhelm the behavior of fermionic quasiparticles, making it exceptionally difficult to analyze electron behavior. For the AE KSDFT calculations, we applied $18\times 18\times 11$ and $18\times 18\times 18$ $k$-grid meshes for the ground-state calculations of Na-hP4 and Al-FCC/Diamond, respectively. The potential and electron density are represented in the dense real-space grids of $120\times 120\times 120$. The eight-order finite-difference method[45] is adopted for evaluating vW kinetic potential.